For the dinner special, one must choose one of seven main entrees, two of four vegetables, and one of three kinds of soup. How many different dinners as described are possible?
Answer by stanbon(75887)

For the dinner special, one must choose one of seven main entrees, two of four vegetables, and one of three kinds of soup. How many different dinners as described are possible?
-----
# of different dinners = 7*4C2*3 = 21*6 = 126
=================
Cheers,
Stan H.
=================